OpenStudy (anonymous):

If a line contains the point (0, -1) and has a slope of 2, then which of the following points also lies on the line?

OpenStudy (calculusxy):

you need to get the slope-int form first \(y = mx + b\) you already have the slope \(m = 2\) and the x and y values (0, -1). find \(b\) \(\color{red}{-1 = 2(0) + b}\) solve for b and then u should get the slope-int form.

OpenStudy (anonymous):

-1?

OpenStudy (calculusxy):

correct :)

OpenStudy (calculusxy):

so we now have: \(y = 2x -1\)

OpenStudy (anonymous):

is it (2,1) ?

OpenStudy (calculusxy):

plug in the x and y value... \(1 = 2(2) - 1\) \(1 = 4 - 1\) \(1 \neq 3\)

OpenStudy (calculusxy):

(2,1) does not fit in the line

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is it (1,1)?

OpenStudy (calculusxy):

\(1 = 2(1) - 1\) \(1 = 2 - 1\) \(1 = 1\) <-- that's correct :)

OpenStudy (anonymous):

ok thank you so much :)

OpenStudy (calculusxy):

yw